OpenStudy (anonymous):

2x+6y=-8 x-3y=8 i think x=8, and y=-4, but i need clarification on what to do.

OpenStudy (whpalmer4):

Okay, first thing I would suggest is checking your solution in both of those equations. How did you arrive at it?

OpenStudy (anonymous):

ok. 2x+6y=-8 x-3y=8 multiplied the bottom equation by two, 2x+6y=-8 2x-6y=16

OpenStudy (whpalmer4):

Okay, then what did you do?

OpenStudy (anonymous):

2x=16/2=(x=8) 2(8) +6y=-8 16+6y=-8

OpenStudy (anonymous):

6y must equal 4. 6(4)=24 16+24=-8

OpenStudy (whpalmer4):

No, you need to add the equations together. \[2x + 6y = -8\]\[2x-6y=16\] --------------------- \[4x + 6y - 6y = -8 + 16\] Now solve it.

OpenStudy (anonymous):

4x=-8/4, x=2

OpenStudy (whpalmer4):

There you go. Now use that plus either of the original equations to find y.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

wait...how did you get 2x if you add it?

OpenStudy (whpalmer4):

What do you get for y? And does your solution make both equations work? I added 2x + 2x = 4x. I also added 6y + (-6y) = 0y. On the right side, I added -8 + 16 = 8. Altogether that gives 4x = 8, x = 2.

OpenStudy (anonymous):

yeah. so you multiply then add. ok. y equals 2.

OpenStudy (anonymous):

sorry, -2.

OpenStudy (whpalmer4):

\[2(2) + 6(-2) = -8\checkmark\]\[2(2) - 6(-2) = 16\checkmark\] Always good to check the solutions!